OpenStudy (elisuzsmith):

The table shows the number of minutes a member of the basketball team played and the points she scored in each of nine games. What is the equation of a trend line that models a relationship between the number of minutes and the number of points? minutes played(x): 18, 17, 10, 28, 18, 23, 30, 16, 24 points scored (y): 15, 10, 6, 18, 12, 14, 16, 3, 10

OpenStudy (anonymous):

the equation of a line is y = mx + c , where m is the slope and c is contant to find slope(m) use the formula y2 - y1/ x2 - x1 take corresponding values then take 2 corresponding values of x & y and use th slope to find the constant.

OpenStudy (anonymous):

for example , x 2 , 4 ,6 y 5 , 8 , 10 y2 = 8 , x2 = 4 & y1 = 5 & x1 =2 so 8-5 / 4-2 = 3/2 is slope

OpenStudy (anonymous):

tell me if u need help

OpenStudy (elisuzsmith):

Thanks!